# Unveiling Downstream Performance Scaling of LLMs: A Clustering-Based Perspective

### v1

@@ 6 Conclusion and Discussion
In this work, we introduce a novel downstream performance scaling framework including (1) a difficulty-based clustering approach that effectively models the underlying distribution of each evaluation set; (2) a theoretically grounded scaling law for downstream task performance that provides a fitting formula for performance-compute extrapolation; and (3) a systematic methodology for identifying and leveraging predictable subset that provides a robust intermediate metric for accurate full-set performance predictions.
Our framework, while effective for dense transformers, has not been fully explored for cost-efficient MoE models and does not account for the annealing phase in training, where high-quality data can rapidly enhance performance. The COD method requires sufficient test cases and is not suited for multiple-choice tasks, where performance metrics may diverge from true passrates. Additionally, the framework’s theoretical foundation is insufficient for chain-of-thought reasoning, necessitating future adaptations to address these challenges. We provide detailed discussion in Appendix E.
Looking forward, our approach can be further expanded across model architectures, training methods, and evaluation set types, while extending this framework to address chain-of-thought reasoning patterns offer promising avenues for future research.
@@ A.1 Improved MeanShift Algorithm
Filtering Zero-performance Samples
@@
Therefore, we pre-filter these zero-performance samples before clustering, treating them as outliers that do not participate in the clustering process. This approach eliminates the need to consider their metrics under large models during subsequent extrapolation and prevents disruption to the clustering of normal difficult samples.
@@ A.2 Smoothing Techniques
Horizontal smoothing: adjacent checkpoint smoothing. Metric fluctuations of individual samples in downstream tasks are not solely due to limited sampling. Another potential factor is noise from uneven data distribution in recent training batches. Therefore, in addition to performing 100 evaluations to mitigate sampling variance, we evaluated 100 times on each of the adjacent checkpoints before and after the selected model. We then averaged these accuracy expectation values across three checkpoints, further reducing sampling variance while offsetting noise from uneven training data distribution. This approach also reduces the number of zero-performance samples, further improving clustering and prediction effectiveness.
Vertical smoothing. Each sample’s features represent the expected correct response rate across models of increasing size, forming a partially ordered sequence. However, the Euclidean distance used for measurement does not consider this sequential information. For example, if a cluster center has a feature sequence of $[0,0,0,0.5]$, sample A with $[0,0,0.2,0.5]$ and sample B with $[0.2,0,0,0.5]$, sample A clearly fits the cluster better than sample B, yet their Euclidean distances are identical.
Note that this smoothing method may not be effective for all downstream tasks. Our current observations suggest that for freeform tasks with limited solution spaces (such as multiple choice, ordering, or judgment questions in freeform format), once models learn to answer within the solution space, their random guess metrics on the evaluation set will be non-zero, more significantly affected by recent training batch data and few-shot cases in prompts. In such cases, vertical smoothing is more likely to bring positive benefits.
In our experiment, we only adopt horizontal smoothing, and leave vertical smoothing as an optional selection.
Appendix B Proof of Proposition
We use Proposition B.1 to derive scaling law for downstream task performance (Proposition B.2).
Proposition B.1 (Arithmetic-geometric mean difference).
@@ Appendix B Proof of Proposition
$\Delta=\mu_{a}-\mu_{g}=\frac{1}{n}\sum_{i=1}^{n}x_{i}-\left(\prod_{i=1}^{n}x_{i}\right)^{\frac{1}{n}}=\frac{\sigma^{2}}{2\mu_{a}}+o(\mu_{a})$ (4)
@@
$\log x=\log\mu+\frac{x-\mu}{\mu}-\frac{(x-\mu)^{2}}{2\mu^{2}}+o\left((x-\mu)^{2}\right)$ (6)
@@
$\log(GM)$ $=\frac{1}{n}\sum_{i=1}^{n}\log x_{i}$
$=\log\mu+\frac{1}{n}\sum_{i=1}^{n}\left(\frac{(x_{i}-\mu_{a})}{\mu_{a}}-\frac{(x_{i}-\mu_{a})^{2}}{2\mu_{a}^{2}}\right)+o(\mu_{a})$
@@
$\mu_{a}-\mu_{g}=\mu_{a}\left(1-\exp{\left(-\frac{\sigma^{2}}{2\mu^{2}}\right)}\right)+o(\mu_{a})$ (7)
@@
Proposition B.2 (Scaling law for downstream task performance).
@@
    The relationship between the answer loss and compute follows a power law, $\frac{1}{n}\sum_{(q,\mathrm{ans})}\mathrm{loss}_{\mathrm{ans}}(C)\sim aC^{-b}+c$;
@@
    For tasks with a finite answer set, the model gives a random guess choice if it cannot truly solve it;
@@
    Task passrate equals the product of the predicted probability of each token $p_{\mathrm{ans}}=\prod_{t\in\mathrm{ans}}p(t)$, which means that each task has unique answer and model output answer only without thinking progress.
@@
$\mathbb{E}_{p}[\mathrm{Acc}(C)]=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{2\mu}\right)+o(\mu)$ (9)
where:
  • $g$ represents the random guess performance floor;
  • $a,b,c$ are positive constants;
  • $\mu=\frac{1}{\#P}\sum_{(q,\mathrm{ans})\in P}loss_{\mathrm{ans}}$;
  • $\sigma^{2}=\frac{1}{\#P}\sum(\mathrm{loss}_{\mathrm{ans}\_t}-\mu)^{2}$.
@@
We first use assumption 3 to establish the relationship between model passrate and loss on a task.
$-\log(p_{\mathrm{ans}})=-\log\left(\prod_{t\in\mathrm{ans}}p(t)\right)=-\sum_{t\in\mathrm{ans}}\log(p_{t})=\mathrm{loss}_{\mathrm{ans}}$ (10)
@@
$\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]$ $=\mathbb{E}_{p}[\exp(-\mathrm{loss}_{\mathrm{ans}}(C))]$ (11)
$=\frac{1}{n}\sum_{(q,\mathrm{ans})\in P}\exp(-\mathrm{loss}_{\mathrm{ans}}(C)).$ (12)
We can adopt Proposition B.1 to switch from arithmetic mean to geometric mean of $\mathrm{loss}$, and apply the power law assumption 1.
@@
Proposition B.2 demonstrates that a metric of an evaluation set with similar difficulty features can be effectively modeled using the following formula:
$f(C)=g+(1-g)\exp{(-aC^{-b}-c)}$ (19)
Appendix C Additional Ablation Studies
C.1 Comparison of Clustering Methods on Extra Evaluation Sets.
We provide additional clustering evaluation results across more evaluation sets in Table C1 and Table C2, which maintain consistency with the conclusions presented in the main text.
Table C1: Clustering performance on advanced task benchmarks (IAD: Intra-cluster Average Distance, OR: Outlier Rate)
@@ Appendix C Additional Ablation Studies
Table C2: Prediction errors on advanced task benchmarks.
@@ C.2 Interpolation Method
The comparative results across different benchmarks are shown in Table C3. We report the prediction error between the real performance of a large model and the mapping result.
Table C3: Comparison of different interpolation methods across benchmarks.
@@
Furthermore, in Fig. C1, we plot the mapping process using different interpolation formulas, where the x-axis represents the predictable subset indices and the y-axis represents the full set indices. The red points are the numerical values to be fitted, green points represent predicted values, purple points represent anchor points and blue points show the actual performance.
@@
The prediction performance shows certain robustness across different interpolation methods. We aim to use the simplest possible interpolation function while maintaining low prediction errors. Based on the above results, we observe that the Cubic Polynomial shows larger prediction errors due to underfitting on BBH, while the Cubic Spline exhibits some overfitting. Both Quartic Polynomial and Quintic Polynomial perform well, therefore we chose the Quartic Polynomial method as it requires fewer fitting parameters.
@@
Figure C1: Performance mapping with different interpolation methods on BBH evaluation set. The cubic spline is overfitted, and the cubic polynomial method is underfitted. Quartic polynomials and quintic polynomials are comparable while quartic polynomial has fewer parameters.
Appendix D Difficulty Distribution of Predictable Subset
@@
We analyze the proportion of predictable subset tasks across different difficulty levels. The difficulty distributions of predictable subset versus complete sets for different evaluation benchmarks are illustrated in Fig. D2. We use the scores from the 12B model as the basis for difficulty classification. The results show that MMLU-pro and GSM8k evaluation sets have larger proportions of predictable subset, indicating that most questions in these datasets exhibit good performance scaling properties. In contrast, many difficult questions with near-zero scores in the Math evaluation set fall outside the predictable subset, requiring adjustment during the mapping phase. Meanwhile, BBH shows consistent proportions of predictable subset across difficulty levels, as some of its questions demonstrate oscillatory patterns with limited improvement despite increased computing.
@@
The proportion of predictable subset can serve as a metric for assessing evaluation set quality. Evaluation sets with larger predictable subset yield more reliable experimental conclusions from smaller models. When constructing evaluation sets, we recommend screening or supplementing unpredictable clusters and ensuring a minimum number of questions for each difficulty feature to reduce metric volatility.
@@
Figure D2: Difficulty distribution comparison on a 12B model between predictable subset and full evaluation set.
Appendix E Limitations
@@
Influence of model structure and training configurations. Mixture-of-Experts (MoE) models excel in training and inference cost, and are widely used in production. In this work, we reveal the performance scaling on dense transformers, while prediction on MoE models is still underexplored. However, we believe that the proposed method is not significantly affected by the model architecture. If we apply the complete pipeline to MoE models, we expect to achieve similar results. Besides, we only study the pre-training performance prediction with a constant learning rate and do not cover the impact of the annealing training. In this phase, higher-quality data is usually adopted, which can rapidly improve the model’s capabilities. As a result, performance prediction faces greater challenges.
@@
Category of evaluation sets. The proposed Clustering-on-Difficulty method requires a sufficient number of test cases, as too few samples can lead to unstable cluster metrics and ineffective estimation. However, from an evaluation set design perspective, an evaluation set with good predictive properties enables more effective generalization from small-scale to large-scale models, thus providing better guidance for model iteration.
@@
Furthermore, for multiple-choice tasks, the model only needs to assign a higher probability to the correct option compared to others, creating a discrepancy between this metric and the model’s true passrate. Given that more evaluation sets are adopting the Chain-of-Thoughts prompts, we have not included multiple-choice tasks that only require option selection.
@@
Chain-of-thought performance prediction. Proposition B.2 assumes that evaluation sets directly assess models’ ability to provide answers. However, increasingly more evaluations allow models to think before providing answers. Recent works on inference time scaling [31, 4] further demonstrate that for tasks involving mathematics, reasoning, and coding, training models to complete tasks through longer inference computation can significantly improve downstream task performance. In cases where the reasoning process or answers are not unique, the relationship between a model’s answer loss and passrate on a task may not necessarily follow the exponential relationship between the answer loss and the sample passrate. Although our approach maintains its prediction effectiveness in such situations, the theoretical explanation for these cases is insufficient. Therefore, we consider improving prediction methods based on chain-of-thought characteristics and expanding theoretical foundations as future work.

### v2

@@ 6 Conclusion and Discussion
In this work, we introduce a novel downstream performance scaling framework including (1) the COD method that effectively models the underlying diverse scaling patterns of evaluation tasks; (2) a scaling law for downstream task performance that provides a fitting formula for performance-compute extrapolation; and (3) a systematic methodology for identifying and leveraging predictable subset that provides a robust intermediate metric for accurate full-set performance predictions. We discuss the limitations of the COD method in Appendix G. Looking forward, our approach can be further expanded across model architectures, training methods, and evaluation set types, while extending this framework to address chain-of-thought reasoning patterns offer promising avenues for future research.
Boarder Impact
The LLMs performance scaling prediction method proposed in this work can be used to guide decision-making for training LLMs and to monitor whether the LLM training process aligns with expectations. Accurate prediction allows for better planning of compute and human resources, thereby playing a valuable role in scientifically advancing the training of larger LLMs. However, if the prediction results are inaccurate, it may lead to outcomes that do not meet expectations for the invested resources, potentially resulting in resource waste or economic losses.
@@ A.1 Improved MeanShift Algorithm
Filtering zero-performance samples.
@@
Therefore, we pre-filter these zero-performance samples before clustering, treating them as outliers that do not participate in the clustering process. This approach obviates the necessity of considering their metrics under large models during subsequent extrapolation, and prevents disruption to the clustering of the remaining samples.
@@ A.2 Smoothing Techniques
Metric fluctuations of individual samples in downstream tasks are not solely due to limited sampling. Another potential factor is noise from uneven data distribution in recent training batches. Therefore, in addition to performing 100 evaluations to mitigate sampling variance, we evaluated 100 times on each of the adjacent checkpoints before and after the selected model. We then averaged these accuracy expectation values across three checkpoints, further reducing sampling variance while offsetting noise from uneven training data distribution. This approach also reduces the number of zero-performance samples, further improving clustering and prediction effectiveness.
Appendix B Proof of Theorem 1
We use lemma 1 to derive the scaling law for downstream task performance (theorem A1).
Lemma 1 (Arithmetic-geometric mean difference).
@@ Appendix B Proof of Theorem 1
$\Delta$ $=\mu_{a}-\mu_{g}=\frac{1}{n}\sum_{i=1}^{n}x_{i}-\left(\prod_{i=1}^{n}x_{i}\right)^{\frac{1}{n}}=\frac{\sigma^{2}}{2\mu_{a}}+o(\mu_{a}).$ (4)
@@
$\log x=\log\mu+\frac{x-\mu}{\mu}-\frac{(x-\mu)^{2}}{2\mu^{2}}+o\left((x-\mu)^{2}\right).$ (6)
@@
$\frac{1}{n}\sum_{i=1}^{n}\log x_{i}=$ $\log\mu+\frac{1}{n}\sum_{i=1}^{n}\left(\frac{(x_{i}-\mu_{a})}{\mu_{a}}-\frac{(x_{i}-\mu_{a})^{2}}{2\mu_{a}^{2}}\right)+o(\mu_{a})$
@@
$\mu_{a}-\mu_{g}=\mu_{a}\left(1-\exp{\left(-\frac{\sigma^{2}}{2\mu^{2}}\right)}\right)+o(\mu_{a}).$ (7)
@@
Theorem A1 (Scaling Law for Downstream Task Performance).
@@
    Given a task $(q,\mathrm{ans})\in P$, the relationship between the answer loss and the compute follows a power law, $\frac{1}{n}\sum_{(q,\mathrm{ans})}\mathrm{loss}_{\mathrm{ans}}(C)\sim aC^{-b}+c$;
@@
    For tasks with a finite answer set, the model gives a random guess choice if it cannot truly solve it, i.e. $\mathbb{E}[\mathrm{Acc|incorrect}]=g$;
@@
    Task passrate equals the product of the predicted probability of each token $p_{\mathrm{ans}}=\prod_{t\in\mathrm{ans}}p(t)$, which means that each task has a unique answer and model output answer only without thinking progress.
@@
$\mathbb{E}_{p}[\mathrm{Acc}(C)]=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{\mu}\right)+o(\mu),$ (9)
where $g$ represents the random guess performance floor, $1-g$ represents the maximum achievable performance improvement, $a,b,c$ are positive constants, and $\mu=\frac{1}{\#P}\sum_{(q,\mathrm{ans})\in P}\mathrm{loss}_{\mathrm{ans}}$; $\sigma^{2}=\frac{1}{\#P}\sum(\mathrm{loss}_{\mathrm{ans}_{t}}-\mu)^{2}$.
@@
We first use Assumption 3 to establish the relationship between model passrate and loss on a task:
$-\log(p_{\mathrm{ans}})$ $=-\log\left(\prod_{t\in\mathrm{ans}}p(t)\right)=-\sum_{t\in\mathrm{ans}}\log(p_{t})=\mathrm{loss}_{\mathrm{ans}}.$ (10)
@@
$\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]$ $=\mathbb{E}_{p}[\exp(-\mathrm{loss}_{\mathrm{ans}}(C))]=\frac{1}{n}\sum_{(q,\mathrm{ans})\in P}\exp(-\mathrm{loss}_{\mathrm{ans}}(C)).$ (11)
We can adopt lemma 1 to switch from arithmetic mean to geometric mean of $\mathrm{loss}$, and apply the power law assumption 1.
@@
theorem A1 demonstrates that a metric of an evaluation set with similar difficulty features can be effectively modeled using the following formula:
$f(C)=g+(1-g)\exp{(-aC^{-b}-c)}.$ (18)
Appendix C Experiment Settings
Model training. To establish performance predictions for large language models, we conduct systematic experiments with a suite of smaller-scale models across different parameter counts and training data volumes, while controlling for other training configurations such as learning rate, batch size, and additional hyperparameters. All models are trained on a constant learning rate scheduler with a few-step warmup stage and data with the same distribution. We list model configurations in Table A1.
Table A1: Model architecture specifications across different sizes.
122M 238M 411M 652M 973M 1.9B 7B 12B 70B (Target)
Param. (M) 122 238 411 652 973 1,901 6,980 12,022 68,452
Tokens (B) 26 45 72 108 153 277 923 1,544 8,012
Continue-Trained Tokens (B) 3 5 8 12 18 33 114 191 1,000
Model Dimension 1,024 1,280 1,536 1,792 2,048 2,560 4,096 4,608 8,192
Heads 8 10 12 14 16 20 32 36 64
KV Heads 8 10 12 14 16 20 8 12 8
Evaluation settings. We conducted performance scaling estimation experiments across 8 major LLM evaluation sets. These evaluation sets span a diverse range of capabilities, including Math, Reasoning, Knowledge, Coding, Reading, and general abilities. All pretrained LLMs were evaluated using a few-shot methodology to obtain the performance metrics. Detailed information is provided in Table A2.
Table A2: Information of evaluation datasets.
Dataset Domain #Questions #Shots
GSM8K Math 1,319 8
MATH Math 5,000 4
BBH Reasoning 6,511 3
TriviaQA Knowledge 17,944 5
MBPP Coding 500 3
AGIEval Comprehensive 8,063 5
DROP Reading 9,536 3
MMLU-pro Comprehensive 12,032 5
Appendix D Performance Prediction for Continue-Pretrained LLMs
Leading industry pre-trained LLMs (e.g., Deepseek-v3 [25], Llama3 [14], Qwen-2.5 [39]) adopt the Continual Pre-Training (CPT) strategy of concentrating high-quality data towards the end of the pre-training process. This phase is typically accompanied by learning rate decay, enabling the model to fully absorb this high-quality data. Due to significant changes in data distribution and the learning rate schedule, this approach often yields substantial improvements in metrics. Predicting a large model’s final capability based solely on its performance during a “stable” phase with consistent data distribution does not reflect its ultimate capability. Therefore, we supplement this by providing metric predictions for the high-quality CPT phase.
The relationship between model parameter scale and the volume of CPT tokens is listed in Table A1.
We conduct the same COD pipeline for CPT models. We control the data distribution of the stable and decay phases for various smaller models, as well as their token-to-parameter ratio, to be consistent with the large model targeted for prediction. The last checkpoint is used for evaluation. Based on prior clustering labels, we perform fitting, extrapolation, and mapping to obtain the predicted performance for the large model.
Table A3: Predicted vs. actual metrics for an LLM with 70B parameters after high-quality continued pretraining.
Evaluation Set Predicted Metric Actual Metric Prediction Error
MMLU-pro 61.58 59.34 2.24
BBH 84.66 85.32 0.66
TriviaQA 79.64 84.05 4.41
MBPP 73.00 73.20 0.20
AGIEval 61.66 64.18 2.52
DROP 81.65 81.39 0.26
gsm8k 86.88 91.81 4.93
MATH 49.29 52.68 3.39
Results listed in Table A3 show that the proposed COD method achieves an average prediction error of 2.33%.
We observe that MATH, GSM8k, and TriviaQA exhibit relatively large prediction errors. We hypothesize that there are two main categories of reasons for this inaccuracy:
  1. 1.
    The CPT data and the evaluation sets possess a significant correlation. For example, in math-related evaluation sets, a modest amount of training can yield substantial improvements in performance metrics. In such scenarios, the metrics for smaller models tend to show greater volatility and have inaccurate evaluations.
  2. 2.
    The CPT data exhibits inherent distribution bias, such that certain evaluation sets, such as TriviaQA, do not derive performance gains from it. This leads to potential significant fluctuations in the metrics after the CPT phase, thereby diminishing the accuracy of extrapolating to larger models.
Appendix E Additional Ablation Studies
E.1 Comparison of Clustering Methods on Extra Evaluation Sets.
We provide additional clustering evaluation results across the remaining evaluation sets in Table A4 and Table A5, which maintain consistency with the conclusions presented in the main text.
Table A4: Clustering performance on advanced task benchmarks (IAD: Intra-cluster Average Distance, OR: Outlier Rate)
@@ Appendix E Additional Ablation Studies
Table A5: Prediction errors on advanced task benchmarks.
@@ E.2 Interpolation Method
The comparative results across different benchmarks are shown in Table A6. We report the prediction error between the real performance of a large model and the mapping result.
Table A6: Comparison of different interpolation methods across benchmarks.
@@
Furthermore, in Fig. A1, we plot the mapping process using different interpolation formulas, where the x-axis represents the predictable subset indices and the y-axis represents the full set indices. The red points are the numerical values to be fitted, green points represent predicted values, purple points represent anchor points, and blue points show the actual performance.
@@
The prediction performance shows certain robustness across different interpolation methods. We aim to use the simplest possible interpolation function while maintaining low prediction errors. Based on the above results, we observe that the Cubic Polynomial shows larger prediction errors due to underfitting on BBH, while the Cubic Spline exhibits some overfitting. Both Quartic Polynomial and Quintic Polynomial perform well, therefore, we chose the Quartic Polynomial method as it requires fewer fitting parameters.
@@
Figure A1: Performance mapping with different interpolation methods on the BBH evaluation set. The cubic spline is overfitted, and the cubic polynomial method is underfitted. Quartic polynomials and quintic polynomials are comparable, while a quartic polynomial has fewer parameters.
Appendix F Difficulty Distribution of Predictable Subset
@@
We analyze the proportion of predictable subset tasks across different difficulty levels. The difficulty distributions of predictable subset versus complete sets for different evaluation benchmarks are illustrated in Fig. A2. We use the scores from the 12B model as the basis for difficulty classification. The results show that MMLU-pro and GSM8k evaluation sets have larger proportions of predictable subsets, indicating that most questions in these datasets exhibit good performance scaling properties. In contrast, many difficult questions with near-zero scores in the MATH evaluation set fall outside the predictable subset, requiring adjustment during the mapping phase. Meanwhile, BBH exhibits consistent proportions of its predictable subset across varying difficulty levels, as some questions display oscillatory patterns with limited improvement, even with increased computational resources.
@@
The proportion of the predictable subset can serve as a metric for assessing evaluation set quality. Evaluation sets with larger predictable subsets yield more reliable experimental conclusions from smaller models. When constructing evaluation sets, we recommend screening or supplementing unpredictable clusters and ensuring a minimum number of questions for each difficulty feature to reduce metric volatility.
@@
Figure A2: Difficulty distribution comparison on a 12B model between predictable subset and full evaluation set.
Appendix G Limitations
@@
Influence of model structure. Mixture-of-Experts (MoE) models excel in training and inference cost, and are widely used in production. In this work, we reveal the performance scaling on dense transformers, while prediction on MoE models is still underexplored. However, we believe that the proposed method is not significantly affected by the model architecture. We expect similar results for MoE models.
@@
Category of evaluation sets. The proposed Clustering-on-Difficulty method requires a sufficient number of test cases, as too few samples can lead to unstable cluster metrics and ineffective estimation. From an evaluation set design perspective, an evaluation set with good predictive properties enables more effective generalization from small-scale to large-scale models, thus providing better guidance for model iteration.
@@
Furthermore, for multiple-choice tasks, the model only needs to assign a higher probability to the correct option compared to others, creating a discrepancy between the answer loss and the model’s true passrate. Given that more evaluation sets are adopting Chain-of-Thoughts prompts, we have not included multiple-choice tasks that only require option selection.
@@
Chain-of-thought performance prediction. theorem A1 assumes that evaluation sets directly assess models’ ability to provide answers. However, increasingly more evaluations allow models to think before providing answers. Recent works on inference time scaling [32, 4] further demonstrate that for tasks involving mathematics, reasoning, and coding, training models to complete tasks through longer inference computation can significantly improve downstream task performance. In cases where the reasoning process or answers are not unique, the relationship between a model’s answer loss and passrate on a task may not necessarily follow the exponential relationship between the answer loss and the sample passrate. Although our approach maintains its prediction effectiveness in such situations, the theoretical explanation for these cases is insufficient. Therefore, we consider improving prediction methods based on chain-of-thought characteristics and expanding theoretical foundations as future work.